(x-1)^2 (x+1)^3
Find the maximum and minimum values
1
Expert's answer
2020-09-22T16:26:08-0400

"\mathbf{Let\;f(x)=(x-1)^2(x+1)^3,\;then\;derivative \;of\;f(x)\;is\;given\;by-}"


"\mathbf{\dfrac{d}{dx}f(x)=f'(x)=\dfrac{d}{dx}\big((x-1)^2(x+1)^3\big)}"


"\mathbf{=(x-1)^2\dfrac{d}{dx}\big((x+1)^3\big)+(x+1)^3\dfrac{d}{dx}\big((x-1)^2\big)}"


"\mathbf{=(x-1)^2.3.(x+1)^{3-1}+(x+1)^3.2.(x-1)^{2-1}}"


"\mathbf{=3(x-1)^2(x+1)^2+2(x+1)^3(x-1)}"


"=\mathbf{(x-1)(x+1)^2[3(x-1)+2(x+1)]}"


"\mathbf{=(x-1)(x+1)^2[5x-1]}"


"\textbf{Now, the derivative (slope) is zero when f'(x)=0 ,i.e.,}"

"\textbf{when x=1,\;x=-1,\;x= 1/5}"


"\textbf{We\;know\;that\;-}"


Second Derivative Test-


When a function's slope is zero at x, and the second derivative at x is:


-less than 0, it is a local maximum

-greater than 0, it is a local minimum

-equal to 0, then the test fails (there may be other ways of finding out though)


Here the second derivative of f(x) is-


"\mathbf{f''(x)=4(5x^3+3x^2-3x-1)}"


"\mathbf{At \;x=1\;, f''(1)=4(5+3-3-1)=4*4=16>0}"

"\mathbf{\implies there\;is\;a\;local\;minimum\;at\;x=1\;and\;at\;x=1\;minimum}"

"\mathbf{value\;of\;the\;function\;is-f(1)=(1-1)^2(1+1)^3=0.}"


"\mathbf{At \;x=-1\;, f''(-1)=4(-5+3+3-1)=4*0=0}"

"\mathbf{\implies the\;test\;fails\;at\;x=-1}"

"\mathbf{\implies there\;is\;a\;saddle\;point\;at\;x=-1,i.e.}"

"\mathbf{\implies there\;is\;neither\;a\;maximum\;nor\;a\;minimum\;at\;x=-1.}"


"\mathbf{At \;x=1/5\;, f''(1/5)=4(5(1/5)^3+3(1/5)^2-3(1/5)-1)}"

"\mathbf{=4*(-36/25)=(-144/25)<0}"

"\mathbf{\implies there\;is\;a\;local\;maximum\;at\;x=1/5\;and\;at\;x=1/5\;maximum}"

"\mathbf{value\;of\;the\;function\;is-f(1/5)=((1/5)-1)^2((1/5)+1)^3}"

"\mathbf{=3456/3125}"